Grade 11IIT JEE Entrance Exam

An aeroplane is flying in the horizontal direction with a velocity 600 km/hr at a height 1960 m. When it is vertically above a point A on the ground, a body is droppd from it. The body strikes a ground at point B. The distance AB is equal to?

Piyanshi Agarwal
7 Years agoGrade 11

1 Answer

Susmita
7 Years ago
The body has a horizontal component of velocity ux=600km/hr and no vertical component uy=0.
Also it has no horizontal component of acceleration ax=0 and ay=g.
The horizontal distance x=uxt.We have to find out t.
h=uyt+gt2/2=gt2/2
or,t2=2h/g
or,t=20s=1/180 hr
So x=uxt=600/180=10/3 km=3.33km
So the distance AB
=root(h2+x2)
=root(19.62+3.332)
=19.8 km
Hope this helps.